# Conditional Generative Modeling for Enhanced Credit Risk Management in Supply Chain Finance

### v1

@@ 3.3.3 Asymptotic Analysis for $r_{3}(l)$ Estimation
This theorem confirms the uniform consistency of the QRGMM estimator for $r_{3}(l)$. The convergence guarantee is facilitated by the boundedness and smoothness of the integrand $g_{l}(\cdot)$, and the controlled behavior of the truncation threshold $a(l)$. These structural conditions are naturally satisfied in SCF applications,
The uniform convergence of $\hat{r}_{3}(l)$, matching the asymptotic behavior of $\hat{r}_{1}(l)$ and $\hat{r}_{2}(l)$, underscores the robustness of QRGMM in the SCF context. This theorem, coupled with the flexibility of $g_{l}(\cdot)$ and $a(l)$, enhancing the applicability of $r_{3}(l)$, allows $r_{3}(l)$ to adapt to diverse risk management needs while maintaining estimation consistency. In summary, the uniform convergence results of $\hat{r}_{1}(l)$, $\hat{r}_{2}(l)$, and $\hat{r}_{3}(l)$ ensures reliable risk assessment across varying loan levels, enhancing decision-making in SCF applications such as adaptive loan sizing and risk mitigation.
@@ 4 Extension of QRGMM for Practical Applicability in SCF Risk Management
In the previous chapters, we focused on the QRGMM under relatively simple model assumptions, notably linear quantile regression. While these assumptions facilitated theoretical derivations and provided valuable insights into convergence properties, they may be restrictive for real-world SCF scenarios. In practice, the relationships governing sales—especially in cross-border e-commerce—are often complex, involving intricate interactions among products, stores, and other contextual factors. Linear models may struggle to capture such combinational effects, potentially undermining the accuracy and effectiveness of distributional predictions required for robust SCF risk management.
To overcome these limitations, this chapter integrates advanced techniques from recommendation systems, specifically Deep Factorization Machines (DeepFM, Guo et al. 2017), with the QRGMM framework. This integration allows us to model non-linear, intricate feature interactions and improve the realism and accuracy of generated sales distributions. We term this extended model as QRGMM, highlighting its advanced capability to handle complex combinational features and non-linear relationships, essential for robust SCF applications.
@@ 4.1 Overview of Related Techniques
This subsection introduces the key building blocks for our extended model. We first discuss Matrix Factorization (MF, Lee and Seung 2000) and Factorization Machines (FM, Rendle 2010), fundamental techniques in recommendation systems that enable efficient modeling of latent and combinational features. We then present Deep Factorization Machines, which combine FM with deep learning technique to capture second-order and higher-order feature interactions. Lastly, the Quantile Regression Neural Network (QRNN, Cannon 2011) provides a remarkable approach for extending quantile regression from linear models to neural networks, making it an ideal glue to seamlessly integrate DeepFM with the QRGMM framework.
@@ 4.1.1 Matrix Factorization (MF)
$\min_{U,V}\sum_{(i,j)\in\mathcal{K}}(r_{ij}-U_{i}^{\top}V_{j})^{2}+\lambda(\|U\|^{2}+\|V\|^{2}),$
@@
In the SCF context, we can analogize users and items to entities such as stores and products, and ratings to observed sales. By identifying latent factors underlying these interactions, MF provides insight into complex, hidden relationships that simpler linear models may fail to capture.
@@ 4.1.2 Factorization Machines (FM)
While MF focuses on a user-item dyad, Factorization Machines extend this idea to a broader class of features and their interactions. FM efficiently captures pairwise feature interactions by introducing latent vectors for each feature:
$\hat{y}(\mathbf{x})=w_{0}+\sum_{i=1}^{n}w_{i}x_{i}+\sum_{i=1}^{n}\sum_{j=i+1}^{n}\langle\mathbf{v}_{i},\mathbf{v}_{j}\rangle x_{i}x_{j},$
[…]
@@
As shown in Figure 2, compared to MF, FM provides a more flexible structure that can incorporate a broader range of features, making it suitable for capturing complex relationships in diverse datasets. Additionally, compared to traditional polynomial kernel-based methods like Support Vector Machines (SVM), FM demonstrates superior generalization capabilities in sparse settings. The dense parametrization of SVMs relies on direct observations of feature interactions, which are often unavailable in sparse data, whereas FM can effectively estimate parameters even under such conditions (for more details, see Rendle (2010)).
@@ 4.1.3 Deep Factorization Machines (DeepFM)
Deep Factorization Machines combine the strengths of FM with deep neural networks to model both low-order and high-order feature interactions. The FM component efficiently learns second-order feature interactions, while the deep neural network component captures complex, higher-order interactions without the need for extensive manual feature engineering. The structure of DeepFM is shown in Figure 3.
@@ 4.1.4 Quantile Regression Neural Network (QRNN)
The methods discussed earlier, including MF, FM, and DeepFM, are primarily designed for expectation-based prediction tasks. However, our goal is to predict the entire distribution of sales, rather than just the mean, which is crucial for robust SCF risk management. Given that DeepFM is essentially a neural network architecture, it becomes necessary to extend the linear quantile regression model of QRGMM into a neural network based quantile regression model. QRNN provides an ideal tool for achieving this extension.
QRNN enhances conventional quantile regression by leveraging neural networks to approximate quantile functions directly. Instead of constraining the quantile function to a linear form, QRNN employs a flexible neural network architecture—often composed of multiple dense layers with nonlinear activation functions—to model the potentially intricate relationship between covariates and various quantiles of the response variable. Unlike the loss functions used in MF, FM, or DeepFM—which typically minimize mean squared error—QRNN adopts the same pinball loss as the linear quantile regression model introduced earlier in this paper. In addition, QRNN provides an efficient way to fit multiple quantiles at different quantile levels $\tau_{j}$, the total loss is typically the sum over all quantiles and samples:
$\min_{\text{model parameters}}\sum_{j=1}^{m}\sum_{i=1}^{n}\rho_{\tau_{j}}(y_{i}-\hat{y}_{i,j}),$
where $\hat{y}_{i,j}$ denotes the predicted $\tau_{j}$-quantile for sample $i$, obtained as the output from the neural network model. The model can be efficiently solved by SGD methods.
By integrating this quantile-specific loss with a neural network structure, QRNN can learn multiple quantiles simultaneously, capturing distinct aspects of the underlying conditional distribution. Neural network components such as nonlinear activations, dropout layers, and proper regularization enable the QRNN to represent complex, non-linear patterns and maintain stability during training. In essence, QRNN extends quantile regression capabilities beyond linear assumptions, allowing for highly non-linear relationships and more realistic modeling of distributions.
@@ 4.2 Integration of DeepFM and QRGMM: QRGMM
Figure 4: Integration of DeepFM and QRNN (QRGMM).
By using DeepFM as the neural backbone of the quantile regression network, we ensure that the learned quantile functions are sensitive to intricate, second-order and higher-order feature interactions. Once the quantile functions are estimated, QRGMM’s inverse transformation approach is applied to generate samples. Specifically, for a given $\mathbf{x}$, we draw $u_{k}\sim\text{Uniform}(0,1)$, and interpolate between the estimated quantiles to generate synthetic observations $\hat{Y}_{k}(\mathbf{x})$ that approximate the conditional distribution of $Y(\mathbf{x})$. The overall procedure follows Algorithm 1, except that the quantile regression model is replaced with the QRGMM model based on DeepFM.
By integrating DeepFM and QRNN within the QRGMM framework, the resulting QRGMM model more accurately captures the complex, non-linear interactions among features that characterize SCF sales patterns, leading to more realistic sales distributions and better estimates of tail risks and other key distributional characteristics. This enhanced generative capability supports more informed lending decisions, improved estimation of risk measures, and stronger risk management strategies. By bridging the gap between linear QRGMM and real-world complexities, QRGMM significantly enhances the applicability of risk management techniques, enabling financial institutions to navigate the uncertainties of cross-border e-commerce markets with greater confidence and more effective risk mitigation.
@@ 5 Numerical Experiments
This section presents a comprehensive evaluation of our proposed QRGMM approach to assess its performance and practical value in SCF risk management. We begin with a synthetic scenario where the full data-generating process is known. In this controlled environment, we can thoroughly examine the model’s ability to accurately estimate conditional quantiles, generate realistic distributions, and precisely approximate various risk measures. Following the synthetic analysis, we apply the method to a real-world datasets, thereby demonstrating its relevance and adaptability to practical SCF contexts.
@@ 5.1 Experiments on Synthetic Data
To create a challenging yet interpretable test environment, we construct a synthetic dataset of 150,000 samples using an FM-Location-Scale Shift model:
$Y(\mathbf{x})=$ $w_{0}+\sum_{i=1}^{n}w_{i}x_{i}+\sum_{i=1}^{n}\sum_{j=i+1}^{n}\left\langle\mathbf{v}_{i},\mathbf{v}_{j}\right\rangle x_{i}x_{j}$
$+\left(r_{0}+\sum_{i=1}^{n}r_{i}x_{i}+\sum_{i=1}^{n}\sum_{j=i+1}^{n}\left\langle\mathbf{z}_{i},\mathbf{z}_{j}\right\rangle x_{i}x_{j}\right)u,$
@@
We randomly partition the data into a training set of 120,000 samples and a test set of 30,000 samples. The number of grid points for the model is set to $m=400$ in accordance with the recommendations provided in Hong et al. (2023) that optimal choice of $m=O(\sqrt{n})$ in practice. After the model is fitted, we assess its performance in terms of quantile estimation accuracy, generative quality, and risk measure estimation through subsequent evaluations, comparing it with the prior popular generative model, Conditional Wasserstein Generative Adversarial Networks with Gradient Penalty (CWGAN-GP, Athey et al. 2021, hereafter referred to as CWGAN).

### v3

@@ 4.2 Asymptotic Analysis for Estimators of Risk Measures
Theorem 1 establishes the uniform consistency of the QRGMM estimator of $r_{3}(l)$. The convergence analysis is facilitated by the boundedness and smoothness of the integrand $g_{l}(\cdot)$, and the controlled behavior of the truncation threshold $a(l)$. These structural conditions are naturally satisfied in the SCF context. The uniform convergence of $\hat{r}_{3}(l)$ and using different risk measures provides a comprehensive basis for reliable risk measures estimation across varying loan levels, supporting adaptive loan sizing.
4.3 Practical Choice of $m$
In the above subsections, we establish the uniform convergence of the estimated conditional distribution function and the estimated risk measures when using the QRGMM algorithm. For implementation, however, one must specify the number of quantile grid points $m$ relative to the sample size $n$. In what follows, we formalize the convergence rate on the main quantile region and translate it into a principled recommendation for $m$.
We work on the main quantile region $\tau_{\ell},\tau_{u}]\subset(0,1)$, where the conditional density $f_{Y}(\cdot\,|\,\mathbf{x})$ is bounded and bounded away from zero by Assumption 2 . Let $Q(\tau\,|\,\mathbf{x})=F_{Y}^{-1}(\tau\,|\,\mathbf{x})$ and $\hat{Q}(\tau\,|\,\mathbf{x})$ be the conditional quantile function and its QRGMM estimator. hong2023learning proves the following convergence rate for the quantile curve on the main quantile region.
Lemma 2 (Hong et.al, 2023).
Under Assumptions 13, for any given $\mathbf{x}$, as $n,m\to\infty$, we have
$\max_{\tau_{j}\in[\tau_{\ell},\tau_{u}]}\bigl|\hat{Q}(\tau_{j}\,|\,\mathbf{x})-Q(\tau_{j}\,|\,\mathbf{x})\bigr|\;=\;O_{\mathbb{P}}\!\left(\frac{1}{\sqrt{n}}\right).$
Moreover,
$\sup_{\tau\in[\tau_{\ell}+1/m,\tau_{u}-1/m]}\bigl|\hat{Q}(\tau\,|\,\mathbf{x})-Q(\tau\,|\,\mathbf{x})\bigr|\;=\;O_{\mathbb{P}}\!\left(\frac{1}{\sqrt{n}}\right)\;+\;O\!\left(\frac{1}{m}\right).$
Here, the notation $O_{\mathbb{P}}(\cdot)$ is defined in the following way: for a sequence $Z_{n}$, $Z_{n}=O_{\mathbb{P}}(h_{n})$ means $Z_{n}/h_{n}$ is bounded in probability. The $O_{\mathbb{P}}(n^{-1/2})$ term reflects statistical error bound from estimating the quantile regressions on the grid, while the $O(m^{-1})$ term is the deterministic interpolation error bound between adjacent quantile levels.
We next convert Lemma 2 into a convergence rate for the estimated conditional distribution function on the corresponding main range of $y$. Define $\mathcal{Y}_{m,n}\left(\mathbf{x}\right)=\bigl[\hat{Q}(\tau_{\ell}+1/m\,|\,\mathbf{x}),\,\hat{Q}(\tau_{u}-1/m\,|\,\mathbf{x})\bigr].$
Proposition 1.
Under Assumptions 13, for any given $\mathbf{x}$, as $n,m\to\infty$, we have
$\sup_{y\in\mathcal{Y}_{m,n}(\mathbf{x})}\bigl|F_{\hat{Y}}(y\,|\,\mathbf{x})-F_{Y}(y\,|\,\mathbf{x})\bigr|\;=\;O_{\mathbb{P}}\!\left(\frac{1}{\sqrt{n}}\right)\;+\;O\!\left(\frac{1}{m}\right).$
The proof of Proposition 1 is provided in the Online Appendix A. The convergence rate of Proposition 1 can also transfer to the default probability estimation $\hat{r}_{1}(l)$. Recall $r_{1}(l)=\Pr\{L(l)>0\}=\Pr\{Y(\mathbf{x})<l/r\}=F_{Y}(l/r\,|\,\mathbf{x})$; for loan levels $l$ such that $l/r\in\mathcal{Y}_{m,n}(\mathbf{x})$, Proposition 1 yields
$\sup_{l:\,l/r\in\mathcal{Y}_{m,n}(\mathbf{x})}\bigl|\hat{r}_{1}(l)-r_{1}(l)\bigr|\;=\;O_{\mathbb{P}}\!\left(\frac{1}{\sqrt{n}}\right)\;+\;O\!\left(\frac{1}{m}\right).$
This rate highlights two sources of error: the estimation error $O_{\mathbb{P}}(n^{-1/2})$ and the interpolation error $O(m^{-1})$. To prevent the interpolation term from dominating while keeping computation reasonable, Proposition 1 suggests the rule of thumb $m=O(\sqrt{n})$.
For more general risk measures, Proposition 1 does not apply as directly as it does for $\hat{r}_{1}(l)$. Even so, the inequality (4) together with (3) and the proof of Theorem 1 indicate that the main part of the error of these risk measures is influenced by the same convergence rate $O_{\mathbb{P}}(n^{-1/2})+O(m^{-1})$.
Through a practical example, we test the sensitivity of choosing $m=\sqrt{n}$ in the Online Appendix B. The results show that the choice is robust and performs reasonably well.
Remark 3.
For the convergence rate analysis, we only work on the main quantile region for both theoretical and practical reasons. Theoretically, rate analysis near the extreme quantile ($\tau\to 0,1$) is substantially more involved: it requires additional stronger tail assumptions and specialized extreme value theory. In particular, when the conditional density goes to 0 near the boundaries $\tau\to 0,1$, standard empirical process arguments no longer apply, and current literature provides limited theoretical tools to establish uniform convergence rates in these regions (chernozhukov2005extremal, hong2023learning). Practically, quantile estimates are more stable in the main region, which covers the bulk of the data, whereas reliable estimation of extreme quantiles would require a far larger sample size than what is typically available in SCF applications. Moreover, in our SCF setting, we do not need to consider extreme quantiles of the sales distribution $Y(\mathbf{x})$: when $\tau\to 0$, sales are nearly zero and no loan would be issued; while when $\tau\to 1$, sales reach their maximum potential but we only care about whether they are sufficient to cover the loan $l$, not about the extreme upper tail itself. In other words, since the potential loss is naturally bounded above by $\bar{l}$, and negative losses are not of primary concern in credit risk evaluation, restricting attention to the main quantile region can be regarded as a reasonable simplification.
@@ 5 Extension of QRGMM for Practical Applicability in SCF Risk Management
In the previous section, we focused on the QRGMM under relatively simple model assumption, linear quantile regression model. While this assumption facilitated theoretical derivations and provided valuable insights into convergence properties, they may be restrictive for real-world SCF scenarios. In practice, the relationships governing sales, especially in cross-border e-commerce, are often complex, involving intricate combinatorial interactions among products, stores, and other contextual factors. The linear model may struggle to capture such combinatorial effects, potentially undermining the accuracy and effectiveness of distributional predictions required for robust SCF risk management.
To overcome this limitation, this chapter integrates advanced techniques from recommendation systems, specifically Deep Factorization Machines (DeepFM, guo2017deepfm), with the QRGMM framework. This integration allows us to model non-linear, intricate feature interactions and improve the realism and accuracy of generated sales distributions. We term this extended model as QRGMM, highlighting its advanced capability to handle complex combinatorial features and non-linear relationships, essential for robust SCF applications.
@@ 5.1 Overview of Related Techniques
This subsection introduces the key building blocks for our extended model. We first discuss Matrix Factorization (MF, lee2000algorithms) and Factorization Machines (FM, rendle2010factorization). They are fundamental techniques in recommendation systems that enable efficient modeling of latent and combinatorial features. We then present the Deep Factorization Machines technique (DeepFM, guo2017deepfm), which combines FM with deep learning techniques to capture second-order and higher-order feature interactions. Lastly, the Quantile Regression Neural Network (QRNN, cannon2011quantile) offers an effective way to extend quantile regression from linear models to neural networks, serving as a natural bridge that seamlessly integrates DeepFM with the QRGMM framework.
@@ 5.1.1 Matrix Factorization (MF)
$\min_{U,V}\sum_{(i,j)\in\mathcal{K}}\left(r_{ij}-U_{i}^{\top}V_{j}\right)^{2}+\lambda\left(\|U\|^{2}+\|V\|^{2}\right),$
@@
In the SCF context, we can analogize users and items to entities such as stores and products, and ratings to observed sales. By identifying latent factors underlying these observed sales, MF provides insight into complex, hidden relationships that simpler linear models may fail to capture.
@@ 5.1.2 Factorization Machines (FM)
While MF focuses on a user-item dyad, the Factorization Machines technique extends this idea to a broader class of features and their interactions. FM efficiently captures pairwise feature interactions by introducing latent vectors for each feature:
$\hat{y}(\mathbf{x})=w_{0}+\sum_{i=1}^{p}w_{i}x_{i}+\sum_{i=1}^{p}\sum_{j=i+1}^{p}\langle\mathbf{v}_{i},\mathbf{v}_{j}\rangle x_{i}x_{j},$
where $\mathbf{v}_{i}\in\mathbb{R}^{k}$ is the latent vector for the $i$-th feature and $\langle\cdot,\cdot\rangle$ denotes the inner product. One key insight of FM is that it models the coefficients of interactions between different features as the inner product of their latent vectors, effectively decomposing their interactions through these coefficients.
@@
As shown in Figure 3, compared to MF, FM provides a more flexible structure that can incorporate a broader range of features, making it suitable for capturing complex relationships in diverse datasets. Additionally, compared to traditional polynomial kernel-based methods like Support Vector Machines (SVM), FM demonstrates superior generalization capabilities in sparse settings. The dense parametrization of SVMs relies on direct observations of feature interactions, which are often unavailable in sparse data, whereas FM can effectively estimate parameters even under such conditions (for more details, see rendle2010factorization).
@@ 5.1.3 Deep Factorization Machines (DeepFM)
The Deep Factorization Machines technique combines the strengths of FM with deep neural networks to model both low-order and high-order feature interactions. The FM component efficiently learns second-order feature interactions, and the deep neural network captures higher-order interactions, together reducing the need for extensive manual feature engineering. The structure of DeepFM is shown in Figure 4.
@@ 5.1.4 Quantile Regression Neural Network (QRNN)
The methods discussed earlier, including MF, FM, and DeepFM, are primarily designed for mean-based prediction tasks. However, our goal is to predict the entire distribution of sales, rather than just the mean, which is crucial for robust SCF risk management. Given that DeepFM is essentially a neural network architecture, it becomes necessary to extend the linear quantile regression model of QRGMM into a neural-network-based quantile regression model. QRNN provides an ideal tool for achieving this extension.
QRNN enhances conventional quantile regression by leveraging neural networks to approximate quantile functions directly. Instead of constraining the quantile function to a linear form, QRNN employs a flexible neural network architecture, often composed of multiple dense layers with nonlinear activation functions, to model the potentially intricate relationship between covariates and various quantiles of the response variable. Unlike the loss functions used in MF, FM, or DeepFM, which typically minimize mean squared error, QRNN adopts the same pinball loss as the linear quantile regression model introduced earlier in this paper. In addition, QRNN provides an efficient way to fit multiple quantiles at different quantile levels $\tau_{j}$, the total loss is typically the sum over all quantiles and observations:
$\min_{\text{model parameters}}\sum_{j=1}^{m-1}\sum_{i=1}^{n}\rho_{\tau_{j}}(y_{i}-\hat{y}_{i,j}),$
where $\hat{y}_{i,j}$ denotes the predicted $\tau_{j}$-quantile for observation $i$, obtained as the output from the neural network model. The model can be efficiently solved by SGD methods.
By integrating this quantile-specific loss with a neural network structure, QRNN can learn multiple quantiles simultaneously, capturing distinct aspects of the underlying conditional distribution. Neural network components such as nonlinear activations, dropout layers, and proper regularization enable the QRNN to represent complex, nonlinear patterns and maintain stability during training. In essence, QRNN extends quantile regression capabilities beyond linear assumptions, allowing for highly non-linear relationships and more realistic modeling of distributions.
@@ 5.2 Integration of DeepFM and QRGMM: QRGMM
Figure 5: QRGMM: integration of DeepFM and QRNN.
By using DeepFM as the neural backbone of the quantile regression network, we ensure that the learned quantile functions are sensitive to intricate second-order and higher-order feature interactions. Once the quantile functions are estimated, QRGMM’s inverse transform approach is applied to generate observations. Specifically, for a given $\mathbf{x}$, we draw $u_{k}\sim\text{Uniform}(0,1)$, and interpolate between the estimated quantiles to generate synthetic observations $\hat{Y}_{k}(\mathbf{x})$ that approximate the conditional distribution of $Y(\mathbf{x})$. The overall procedure follows Algorithm 1, except that the quantile regression model is replaced with the QRGMM model based on DeepFM.
By integrating DeepFM and QRNN within the QRGMM framework, the resulting QRGMM model more accurately captures the complex, non-linear interactions among features that characterize SCF sales patterns, bridging the gap between linear QRGMM and real-world complexities and leading to more realistic sales distributions. This enhanced generative capability supports improved estimation of risk measures, more informed lending decisions, and stronger risk management strategies.
Remark 4.
While our framework is motivated by credit risk management in 3PL-led CBEC SCF, its methodological core, constructing a reliable generative model of the underlying distribution and estimating risk measures via Monte Carlo simulation, naturally extends to other financial contexts. For example, the proposed approach can accommodate classical risk measures such as value-at-risk (VaR) and conditional value-at-risk (CVaR), which are widely used in insurance and financial risk management. In principle, once a reliable generative model is available, it can be used to simulate losses under different scenarios to estimate a wide spectrum of financial risk measures.
What makes our setting distinctive, however, is the close linkage between sales distributions and collateral value in the CBEC supply chain finance. Here, the repayment of a loan depends directly on realized sales, making the accurate modeling of $Y(\mathbf{x})$ not only relevant but indispensable. In other SCF contexts, such as domestic trade credit, alternative forms of collateral (e.g., factories, equipments, or real estates) often mitigate the reliance on sales forecasts, reducing the need for such distributional modeling. By contrast, in cross-border e-commerce, the absence of tangible collateral elevates the importance of predictive analytics, which motivates our integration of QRGMM with DeepFM to better capture the intricate combinatorial patterns inherent in e-commerce data. Therefore, while the proposed framework is in principle extensible to other domains, its present formulation is particularly well aligned with the unique challenges of CBEC-SCF.
@@ 6 Numerical Experiments
This section presents a comprehensive evaluation of our proposed QRGMM approach to assess its performance and practical value in SCF risk management. We begin with a synthetic scenario where the full data-generating process is known. In this controlled environment, we can thoroughly examine the model’s ability to accurately estimate conditional quantiles, generate realistic distributions, and precisely estimate various risk measures. Following the synthetic analysis, we apply the method to two real-world datasets, thereby demonstrating its relevance and adaptability to practical SCF contexts.
@@ 6.1 Experiments on Synthetic Data
To create a challenging yet interpretable test environment, we construct a synthetic dataset of 10,000 observations using an FM-location-scale shift model:
$Y(\mathbf{x})=$ $w_{0}+\sum_{i=1}^{p}w_{i}x_{i}+\sum_{i=1}^{p}\sum_{j=i+1}^{p}\left\langle\mathbf{v}_{i},\mathbf{v}_{j}\right\rangle x_{i}x_{j}$
$+\left(r_{0}+\sum_{i=1}^{p}r_{i}x_{i}+\sum_{i=1}^{p}\sum_{j=i+1}^{p}\left\langle\mathbf{z}_{i},\mathbf{z}_{j}\right\rangle x_{i}x_{j}\right)u,$
@@
We randomly partition the data into a training set of 8,000 observations and a test set of 2,000 observations. Throughout all experiments, we follow the guideline in Section 4.3 and set $m=\sqrt{n}$, which yields $m=90$ in this case. In addition, a sensitivity analysis of QRGMM with respect to $m$ is presented in the Online Appendix B, confirming that the choice $m=\sqrt{n}$ is reasonable in practice. After the model is fitted, we assess its performance in terms of quantile estimation accuracy, generative quality, and risk measures estimation through subsequent evaluations, and compare it with a popular generative model, the conditional wasserstein generative adversarial networks with gradient penalty (CWGAN-GP, athey2021using, hereafter referred to as CWGAN).